# Starfield: Demand-Aware Satellite Topology Design for Low-Earth Orbit Mega Constellations

### v1

@@ 5.1. Overview
Starfield is a heuristic algorithm (Appendix C) that achieves near-optimal topology performance by selecting a $\kappa$-max-degree subgraph $G$, which represents the network topology, based on the demand matrix $\Delta$ and the in-range satellite graph $R$. In $R$, the satellite set $S$ forms the nodes, and an edge exists between two satellites if they remain within communication range over a time period $T$. Algorithm 1 presents the Starfield topology generation algorithm. Starfield assigns priorities to candidate links based on a distance metric derived from an abstract vector field (Section 5.2) that captures the influence of individual traffic flows, as shown in Lines 3–10. The metric consists of a dominant component, $f^{\perp}{f^{\perp}}^{T}$, which orients links along the vector field, and a small regularization term, $\epsilon I$, which prevents degeneracy and preserves sensitivity to the original geometric distance in regions where the field magnitude is extremely weak (Line 7). Distances, for priority, can be computed as described in Section 2.3. However, due to the prohibitive computational cost of evaluating the associated integral, we approximate the distance by $|\vec{f^{\perp}}\cdot(\vec{P_{s}}-\vec{P_{s^{\prime}}})|$, where $\vec{P_{s}}$ and $\vec{P_{s^{\prime}}}$ denote the position vectors of satellites $s$ and $s^{\prime}$ forming the link (Line 8).
Algorithm 1. Starfield topology generation algorithm
[…]
@@ 5.3.2. Dealing with the Crown
As shown in Figure 2 (left), a large fraction of traffic geodesics between highly populated cities pass through the North Pole region. However, due to the limited orbital inclination, Phase 1 Starlink satellites do not cover this area. This demand pattern highlights the importance of link orientation near the boundary of satellite coverage, which we refer to as the crown. The absence of satellite coverage in the polar region, which is inherently tied to the constellation’s inclination, leads to suboptimal routing paths in Starfield (Appendix B). As noted in (Alexander and Alexander, 1981) and (De Berg et al., 2008), the optimal path on a Riemannian manifold with boundary—the constellation’s shell—follows a geodesic within the interior of the manifold, but aligns with the boundary curve at points where the geodesic intersects the boundary. Therefore, to achieve near-optimal paths under this constraint, links in the crown area should be oriented parallel to the crown’s line of latitude, rather than forming a jagged pattern, which may arise from the original vector field design (Section 5.2). To this end, we modify the vector field ($\vec{f_{c}}$) to gradually orient links toward lines of latitude as satellites approach the crown:
@@ 6.1.2. Baselines
  • Motif: The Motif-based topology scheme (Bhattacherjee and Singla, 2019) formulates topology selection as a search problem driven by a linear objective function. Their results indicate that most of the discovered motifs are heavily biased toward reducing hop count, largely neglecting stretch factor, and consequently fail to outperform +Grid. This behavior suggests an ineffective formulation of the trade-off objective function, stemming from the direct aggregation of metrics with fundamentally different natures—one being an absolute quantity and the other a normalized factor. This limitation is significant, as stretch factor is directly correlated with propagation delay, a dominant component of end-to-end network latency. Additionally, the motif-based approach constrains the topology search space to uniform motif structures, thereby limiting the range of achievable optimal solutions. Furthermore, search algorithms are slow and exhaustive leading to a prohibitive runtime and limited reproducibility. Therefore, we adopt the +Grid topology—mathematically a motif—as a practical and representative baseline.
@@ 6.2.1. Static Starfield
Although Starfield is a dynamic design applicable at any instant in time, it does not explicitly account for inter-satellite motion. To address this limitation, we introduce a simple static modification to the topology design. Since the dynamic behavior of the constellation arises primarily from inter-orbital satellite connections, we enforce a consistent one-to-one connection pattern (i.e., a bipartite assignment) between satellites in orbits. Specifically, satellite pairs that follow a given pattern—or its corresponding angular variant—are constrained to remain within communication range. Adhering to this pattern guarantees a static terrestrial view of the resulting topology. Formally, if we assign a sequential identifier to each satellite in orbit $O$, denoted by $O_{i}$, the selected pattern $p$ is defined as follows:
@@
where $N_{S}$ denotes the number of satellites per orbit, $D_{f}$ represents the distance function (Section 5.1) for flow $f$, $O$ and $O^{\prime}$ are two desired orbits. Accordingly, for each orbit, the static topology is obtained by selecting, across all feasible patterns, the orbit-to-orbit pattern minimizing Equation 7, and its corresponding angular patterns.
@@
Figure 4 shows the cumulative distribution functions (CDFs) of the stretch factor under different traffic demand patterns for the static Starfield and two baselines, +Grid and Random. Starfield consistently outperforms both baselines across all four demand patterns, achieving up to a 20% reduction in stretch factor relative to +Grid and up to a 50% reduction relative to Random under distance-based demand. The performance gain diminishes as the traffic pattern becomes more uniform. In our simulation, highly populated cities are approximately uniformly distributed over the Earth, making the population-based demand pattern the most directionally uniform. In contrast, distance-based and hotspot traffic induce more concentrated, directional flows. As demand becomes increasingly uniform, Starfield’s advantage decreases, which can be attributed to flows arriving from a broader range of directions, making it infeasible to locally accommodate all directions efficiently.
@@
Figure 5 presents link usage, measured by the proportion of total packets relayed on each link as an indicator of congestion, under different traffic demand patterns for the static Starfield topology and the +Grid and Random baselines. Starfield exhibits higher link congestion across all traffic patterns, characterized by a larger fraction of heavily utilized links. This behavior arises because Starfield intentionally routes traffic along shortest paths to minimize stretch factor, which in turn concentrates traffic on these paths. In contrast, the Random topology demonstrates lower congestion, likely due to the presence of multiple randomized shortest paths that distribute traffic more evenly across the network. Furthermore, more uniform traffic patterns result in reduced congestion overall, primarily due to the naturally balanced distribution of flows.
@@ 6.2.2. Dynamic Starfield
Figure 6 illustrates the CDFs of stretch factor (left) and round-trip time (RTT) (right). Both static and dynamic Starfield outperform +Grid in terms of stretch factor, with dynamic Starfield achieving up to a 15% reduction and static Starfield achieving up to 20% reduction relative to +Grid. However, dynamic Starfield underperforms its static counterpart, which can be attributed to temporal inconsistencies in inter-satellite distances. The figure also shows a modest improvement of approximately $10\,\text{ms}$ in RTT for both static and dynamic Starfield compared to +Grid. Although a larger RTT reduction might be expected from improved propagation delay, this gain is offset by increased transmission and queuing delays caused by higher congestion and longer hop counts, ultimately limiting the end-to-end RTT improvement. With additional experiments under the same settings but with a constellation inclination of $80^{\circ}$, we observe that dynamic Starfield achieves a 30% improvement, while static Starfield attains a 25% improvement relative to +Grid. Similarly, a larger reduction in round-trip time (RTT) of approximately $26\,\text{ms}$ is observed (Appendix B). These results verify the importance of crown size and reorientation discussed in Section 5.3.2.
@@ 6.2.3. Starfield Robustness
6.2.4. Ablation Analysis
Figure 8 presents the CDFs of the stretch factor under different traffic demand patterns for Starfield with $K=10^{6}$, Starfield with $K=10^{7}$, and +Grid. Across all demand patterns, the $K=10^{7}$ case configuration consistently outperforms +Grid, whereas the $K=10^{6}$ case configuration underperforms, in accordance with the design principle that larger values of $K$ favor stretch-factor optimization. Figure 9 shows the CDFs of hop count for the same configurations. In contrast to stretch factor, Starfield with $K=10^{6}$, consistently achieves lower hop counts than $K=10^{7}$ configuration across all demand patterns, aligning with the design principle that smaller values of $K$ favor hop-count minimization. An interesting observation across all traffic patterns is that the stretch-factor trade-off is more favorable under uniform demand distributions, whereas the hop-count trade-off is more pronounced under skewed demand patterns.
@@ A.4. Region Flows
Appendix B Crown Size Impact
Figure 12. CDFs of city-to-city stretch factor for +Grid, dynamic Starfield, and static Starfield under constellation inclinations of $80^{\circ}$ (left) and $53^{\circ}$ (right). Vertical lines indicate the $90^{\text{th}}$ percentile.
To evaluate the impact of crown size on Starfield’s performance, we increase the inclination from the Phase 1 Starlink value of $53^{\circ}$ to an arbitrarily higher $80^{\circ}$, thereby reducing the crown area and and enabling a larger fraction of geodesically optimal paths that are less affected by boundary. Figure 12 presents the CDFs of stretch factor for inclinations of $80^{\circ}$ (left) and $53^{\circ}$ (right). Under the $80^{\circ}$ inclination, both static and dynamic Starfield outperform +Grid in terms of stretch factor, with dynamic Starfield achieving up to a 30% reduction and static Starfield achieving up to a 25% reduction relative to +Grid. These improvements exceed those observed under the $53^{\circ}$ inclination, where dynamic and static Starfield achieve up to 15% and 20% reductions, respectively—an effect that is particularly pronounced for the dynamic variant (Section 6.2.2). At the same time, we observe an overall degradation in stretch factor as the inclination increases. This can be attributed to the reduced satellite density resulting from covering an approximately 44% larger surface area with the same number of satellites. The lower density increases inter-satellite distances, leading to longer links that are less effective at closely approximating geodesic paths, as anticipated by the analysis in Appendix C. Complementing the stretch factor improvements, Figure 13 also shows a larger reduction in round-trip time (RTT) of approximately $26\,\text{ms}$ for both static and dynamic Starfield under the $80^{\circ}$ inclination, compared to the $10\,\text{ms}$ improvement observed under the $53^{\circ}$ inclination.
Figure 13. CDFs of city-to-city round-trip time for +Grid, dynamic Starfield, and static Starfield under constellation inclinations of $80^{\circ}$ (left) and $53^{\circ}$ (right). Vertical lines indicate the $75^{\text{th}}$ percentile.
Appendix C Mathematical Analysis
C.1. Model

### v2

@@ 5.1. Overview
Starfield is an algorithm that achieves near-optimal topology performance by selecting a $\kappa$-max-degree subgraph $G$, which represents the network topology, from a set of satellites $S$ based on the corresponding in-range satellite graph $R$ and demand matrix $\Delta$. Intuitively, it constructs a highway-like topology—prioritizing major traffic corridors—rather than a dense, grid-like layout under limited link budgets. Starfield orients links toward major demand flows by following geodesic curves. Toward this objective, Starfield derives an abstract vector field from the geodesic demands. It then defines a Riemannian metric on the shell manifold to assign a distance to each link. The metric captures the influence of individual demand flows by elongating distances perpendicular to the flow’s direction and contracting distances along the flow. Finally, to achieve near-uniform link coverage in all directions such as in +Grid, Starfield uses the Riemannian metric to choose a primary ISL at each satellite, from which other ISLs are derived. E.g., if $\kappa=4$, each satellite chooses one ISL through the Riemannian metric, with the other ISL computed to be perpendicular to the first ISL.
@@ 5.2. Abstract Demand-based Vector Field
5.3. Starfield Algorithm
Algorithm 1 presents the Starfield topology generation algorithm. Starfield assigns priorities to candidate links based on a distance metric derived from an abstract vector field (Section 5.2) that captures the influence of individual traffic flows, as shown in Lines 3–10. The metric consists of a dominant component, $f^{\perp}{f^{\perp}}^{T}$, which orients links along the vector field, and a small regularization term, $\epsilon I$, which prevents degeneracy and preserves sensitivity to the original geometric distance in regions where the field magnitude is extremely weak (Line 7). Distances, for priority, can be computed as described in Section 2.3. However, due to the prohibitive computational cost of evaluating the associated integral, we approximate the distance by $|\vec{f^{\perp}}\cdot(\vec{P_{s}}-\vec{P_{s^{\prime}}})|$, where $\vec{P_{s}}$ and $\vec{P_{s^{\prime}}}$ denote the position vectors of satellites $s$ and $s^{\prime}$ forming the link (Line 8).
Algorithm 1. Starfield topology generation algorithm
[…]
@@ 5.4.2. Dealing with the Crown
As shown in Figure 2 (left), a large fraction of traffic geodesics between highly populated cities pass through the North Pole region. However, due to the limited orbital inclination, Phase 1 Starlink satellites do not cover this area. This demand pattern highlights the importance of link orientation near the boundary of satellite coverage, which we refer to as the crown. The absence of satellite coverage in the polar region, which is inherently tied to the constellation’s inclination, leads to suboptimal routing paths in Starfield. As noted in (Alexander and Alexander, 1981) and (De Berg et al., 2008), the optimal path on a Riemannian manifold with boundary—the constellation’s shell—follows a geodesic within the interior of the manifold, but aligns with the boundary curve at points where the geodesic intersects the boundary. Therefore, to achieve near-optimal paths under this constraint, links in the crown area should be oriented parallel to the crown’s line of latitude, rather than forming a jagged pattern, which may arise from the original vector field design (Section 5.2). To this end, we modify the vector field ($\vec{f_{c}}$) to gradually orient links toward lines of latitude as satellites approach the crown:
@@
5.4.3. Time Complexity
The complexity consists of two parts: distance computation and subgraph selection. Since vector operations are constant-time, the per-flow link distance computation has complexity $\mathcal{O}(m^{2}|E(R)|)$, where $|E(R)|$ denotes the number of edges in $R$, and $m$ is the number of ground stations. Due to $\zeta$ composition of constant-time vector operations, and on-the-fly minimum distance computation, subgraph selection has complexity $\mathcal{O}(\kappa|S|)$. Therefore, the overall complexity can be combined to $\mathcal{O}(m^{2}|E(R)|+\kappa|S|)$.
5.4.4. Analysis
To simplify the mathematical model, a setting is assumed in which the manifold is a two-dimensional plane equipped with the standard Euclidean metric, and satellites remain stationary over a small time window. In this setting, we derive a lower bound on the stretch for a given traffic demand pattern under all possible topologies. We also show an upper bound on the stretch achieved by the Starfield topology.
Theorem 5.1 (informal).
For any demand, the stretch achieved in Starfield is upper bounded by a factor that depends on the sum of the sine and cosine of the angles between the demand’s geodesic and the vector field along the geodeisc.
For demands that do not have other competing demands nearby, we show that Starfield’s stretch is near-optimal. Refer to Appendix B for the precise mathematical theorems and proofs.
@@ 6.1.2. Baselines
  • Motif: The Motif-based topology scheme (Bhattacherjee and Singla, 2019) formulates topology selection as a search problem driven by a linear objective function. Their results indicate that most of the discovered motifs are heavily biased toward reducing hop count, largely neglecting stretch factor, and consequently fail to outperform +Grid. This behavior suggests an ineffective formulation of the trade-off objective function, stemming from the direct aggregation of metrics with fundamentally different natures—one being an absolute quantity and the other a normalized factor. This limitation is significant, as stretch factor is directly correlated with propagation delay, a dominant component of end-to-end network latency. Additionally, the motif-based approach constrains the topology search space to uniform motif structures, thereby limiting the range of achievable optimal solutions. Nevertheless, we use $\Phi_{1}$, the sum of hop count and stretch factor, as the objective function to optimize the motif search.
@@ 6.2.1. Static Starfield
Maintaining stability over time needs addressing inter-orbital distance variations. In static Starfield, we consider a stable and homogeneous topology in which ISLs between satellite pairs are always in range and, therefore, do not require any reconfiguration over an orbital period. We enforce homogeneity by following the same local ISL connectivity pattern on all satellites within an orbit. We compute the static Starfield topology by selecting for each adjacent orbit pair a feasible connection pattern that minimizes the total induced link distances. Formally, assigning indices $O_{i}$ to satellites in orbit $O$, the pattern $p$ is defined as:
@@
where $N_{S}$ denotes the number of satellites per orbit, $D_{f}$ the flow distance function (Section 5.1), and $O,O^{\prime}$ the target orbits.
@@
Figure 4 shows the cumulative distribution functions (CDFs) of the stretch factor under different traffic demand patterns for the static Starfield and three baselines, +Grid, Motif, and Random. Starfield consistently outperforms both baselines across all four demand patterns, achieving up to a 20% reduction in stretch factor relative to +Grid, up to a 40% reduction relative to Motif, and up to a 50% reduction relative to Random under distance-based demand. The performance gain diminishes as the traffic pattern becomes more uniform. In our simulation, highly populated cities are approximately uniformly distributed over the Earth, making the population-based demand pattern the most directionally uniform. In contrast, distance-based and hotspot traffic induce more concentrated, directional flows. As demand becomes increasingly uniform, Starfield’s advantage decreases, which can be attributed to flows arriving from a broader range of directions, making it infeasible to locally accommodate all directions efficiently.
@@
Figure 5 presents link usage, measured by the proportion of total packets relayed on each link as an indicator of congestion, under different traffic demand patterns for the static Starfield topology, +Grid, Motif, and Random baselines. Starfield exhibits higher link congestion across all traffic patterns, characterized by a larger fraction of heavily utilized links. This behavior arises because Starfield intentionally routes traffic along shortest paths to minimize stretch factor, which in turn concentrates traffic on these paths. In contrast, the Motif topology demonstrates lower congestion, likely due to the presence of multiple randomized shortest paths that distribute traffic more evenly across the network, and lower hop counts. Furthermore, more uniform traffic patterns result in reduced congestion overall, primarily due to the naturally balanced distribution of flows.
@@ 6.2.2. Dynamic Starfield
Figure 6 illustrates the CDFs of stretch factor (left) and round-trip time (RTT) (right). Both static and dynamic Starfield outperform +Grid in terms of stretch factor, with dynamic Starfield achieving up to a 15% reduction and static Starfield achieving up to 20% reduction relative to +Grid. However, dynamic Starfield underperforms its static counterpart, which can be attributed to temporal inconsistencies in inter-satellite distances. The figure also shows a modest improvement of approximately $10\,\text{ms}$ in RTT for both static and dynamic Starfield compared to +Grid. Although a larger RTT reduction might be expected from improved propagation delay, this gain is offset by increased transmission and queuing delays caused by higher congestion and longer hop counts, ultimately limiting the end-to-end RTT improvement.
@@
6.2.3. Crown Size Impact
Figure 8. CDFs of city-to-city stretch (left) and RTT (right) for +Grid, dynamic Starfield, and static Starfield under constellation inclinations of $80^{\circ}$; left lines mark the 90th percentile, right lines the 75th percentile.
To evaluate the impact of crown size on Starfield’s performance, we increase the inclination from the Phase 1 Starlink value of $53^{\circ}$ to an arbitrarily higher $80^{\circ}$, thereby reducing the crown area and and enabling a larger fraction of geodesically optimal paths that are less affected by boundary. Figure 8 presents the CDFs of stretch factor and round trip time for inclination of $80^{\circ}$. Under the $80^{\circ}$ inclination, dynamic Starfield improves performance by 30% and static Starfield by 25% versus +Grid, reducing RTT by $26\,\text{ms}$. These improvements exceed those observed under the $53^{\circ}$ inclination, where dynamic and static Starfield achieve up to 15% and 20% reductions, respectively—an effect that is particularly pronounced for the dynamic variant (Section 6.2.2). At the same time, we observe an overall degradation in stretch factor as the inclination increases. This can be attributed to the reduced satellite density resulting from covering an approximately 44% larger surface area with the same number of satellites. The lower density increases inter-satellite distances, leading to longer links that are less effective at closely approximating geodesic paths, as anticipated by the analysis in Appendix B.
6.2.4. Ablation Analysis
Figure 9 presents the CDFs of the stretch factor under different traffic demand patterns for Starfield with $K=10^{6}$, Starfield with $K=10^{7}$, and +Grid. Across all demand patterns, the $K=10^{7}$ case configuration consistently outperforms +Grid, whereas the $K=10^{6}$ case configuration underperforms, in accordance with the design principle that larger values of $K$ favor stretch-factor optimization. Figure 10 shows the CDFs of hop count for the same configurations. In contrast to stretch factor, Starfield with $K=10^{6}$, consistently achieves lower hop counts than $K=10^{7}$ configuration across all demand patterns against Motif—which overemphasizes hop-count optimization, aligning with the design principle that smaller values of $K$ favor hop-count minimization. Notably, both stretch and hop count gains are more pronounced under uniform demands rather than skewed traffic.
@@ A.4. Region Flows
Appendix B Mathematical Analysis
B.1. Model